The balance sheet provides a snapshot of the financial condition of a company. Investors and analysts use the information given on the balance sheet and other financial statements to make several interpretations regarding the companys financial condition and Puzry Button Clothing Company is a hypothetical company. Suppose it has the following balance sheet tems reported at the end of to complete the balance Liablities and eauty Current habities $6,457 Accounts payable so s 2,363 Aceruals $19,687 $15,750 Total current liabilties 19,250 Total debt $8,750 Common equity 21,328 Retained earnings $43,750 in the preceding balance sheet-and assuming that Fuzzy Button Clothing Company has 50 milion shares of O Type here to search
Chapter 3 Assignment This statement is O Notes payable actually increased from $2,324 million to $2,187 million between Years 1 and 2 Q Accruals actually increased from S0 r, Year 1 to $410 million at the end of Year 2 O Long-term debt decreased from $2,324 million at the end of Year 1 to $2,187 millon by the end of Year Statement 2: In Year 2, Fuzzy Button Clothing Company was profitable This statement is because: O Fuzzy Buttons retained eamings account increased between the end of Years 1 and 2 O Fuzzy Buttons total assets increased between Years 1 and 2 O The cash and equivalents account increased between Years 1 and 2 Statement 3: The book value per share of Fuzzy Buttons stock in Year 2 mas $656.26. because: O The per-share book value is calculated by dividing the companys total debe by the number of 0 The per share book vane is calculated by dividngte companys total assets by the number of O The per-share book value is calculated by dividing the companys total common equity by the number of outstanding shares of common stock shares of common stock outstanding shares of common stock Based on your statement regarding Fuzry Buton Clothing Companys Principles (GAAP) of the different items reported in the balance sheet and the information they provide, which balance sheet is consistent with U.S. Generally Accepted Accounting O The companys assets should be lsted in the order in which they are to be converted into cash. O The companys assets should be lsted in alphabetical order. O The companys assets should be listed from those carrying the largest balance to those with the O Type here to search 0
Answer #1

Please provide rating...

Balance sheet missing figure ..
a Cash and cash equivalent =19687-8662-2953
Total current asset - AR - Inventory 8072
b Net plant and equipment =43750-19687
Total asset - Current asset 24063
c Total Current liabilities =0+410+2324
Account payable+Accruals+Note payable 2734
d Retained earning =32813-21328
Total common equity - Common stock 11485
Statement # 1
Statement is False : Accruals actually increased from 0 in year 1 to $410 million at the end of year 2.
Statement # 2
Statement is True :Fuzzy button's retained earning account increased between the end of year 1 and 2
Statement # 3
Statement is True :the per share book value is calculated by dividing the company's total common equity by the number of outstanding share of common stock.
Q -5 Correct statement is : The company's assets should be listed in the order in which they are to be converted into cash.